Can you please explain how to eshtablish relation among different physical quantities using dimensional analysis?

Grade:10

1 Answers

Aarti Gupta
9 years ago
The dimensional analysis involves use of dimensional formulae of physical quantities which can be obtained from the equations representing their relation with other physical quantities,whose dimensions in mass,length and time are already known.,For ex-
1. Velocity = displacement/ time = [L]/[T]
2. Acceleration = Change in velocity/time taken
a = L/T/T = LT-2 = [M0LT-2]
Similarly,
3. force = mass * acceleration
F = [M] * [LT-2]
F = [MLT-2]
This is how relation among different physical quantities can be established using dimensional analysis.